Circular No. 8477

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2004 MN_{4}

On 2004 Dec. 20, K. E. Smalley (cf. MPEC 2004-Y25) identified an object placed on the ‘NEO Confirmation Page’ on Dec. 18 (following its discovery by G. J. Garradd via the Siding Spring Survey) with 2004 MN_{4} (which had been recorded at Kitt Peak on June 19 and 20; cf. MPS 109613). Although the recognition of further prediscovery observations by Spacewatch on Mar. 15 (MPEC 2004-Y70) precluded the possible 2029 Apr. 13 earth impact discussed extensively in the WWW during 2004 Dec. 23–27 [notably on the Jet Propulsion Laboratory (JPL) and Pisa NEODys “risk pages”], it was clear that the object would then make an unusually close approach. L. A. M. Benner, JPL; M. C. Nolan, National Astronomy and Ionosphere Center, Arecibo Observatory; J. D. Giorgini, S. R. Chesley, and S. J. Ostro, JPL; and D. J. Scheeres, University of Michigan, report: “Arecibo delay-Doppler radar astrometry obtained on 2005 Jan. 27, 29, and 30 significantly refines the 2004 MN_{4} orbit. On Jan. 29.0 UT, the range was 294 km closer to the earth than the pre-radar orbit solution predicted. This correction results in a 2029 approach to the geocenter of only 0.000245 ± 0.000060 AU (36700 ± 9000 km or 5.7 ± 1.4 earth radii, 3σ uncertainties), which is just below geosynchronous orbit and 28000 km closer than predicted by the pre-radar ephemeris. During its close approach, it is likely that tidal torques will significantly alter the object’s spin state.”

ARP 299 Further to IAUC 8473, S. Mattila, R. Greimel, C. Gerardy, and W. P.

S. Meikle, together with D. L. Clements and K. Nandra (Imperial College, London), report the discovery on a K_{s}-band image obtained on Jan. 30.3 UT of a strong outburst in the B1 nucleus (Wynn-Williams et al. 1991, Ap.J. 377, 426) of the galaxy Arp 299. The nucleus was not in outburst in their previous such image taken on 2004 June 6.0, and the difference in brightness yields a magnitude for the new source of K_{s} = +12.6 ± 0.1; at a distance of 45 Mpc, this translates to M(K_{s}) ∼ −20.7, which is about a magnitude brighter than would be expected for the brightest supernovae (cf. Mattila and Meikle 2001, MNRAS 324, 325). The new source is located at α = 11^{h}28^{m}31^{s}.02, δ = +58^{o}33′40′′.7 (equinox 2000.0), which is < 0′′.1 from the K_{s}-band nucleus B1 and < 0′′.5 from the location of the x-ray-determined active galactic nucleus (AGN) in Arp 299 (Ballo et al. 2004, Ap.J. 600, 634). The source is therefore most likely associated with an AGN outburst in the nucleus B1.

2005 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green